OpenStudy (anonymous):

Emily is converting 9 feet per minute into inches per minute. Which conversion factor should she use?

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

omg that's my name :O

ganeshie8 (ganeshie8):

\(\large \frac{9~ft}{1~minute} \times \frac{12 ~in}{1~ft}\)